# Revisiting Schwarzschild black hole singularity through string theory

### v2

@@ Abstract
In this letter, we derive the singular condition for black holes and demonstrate the potential resolution of the black hole singularity in general relativity using non-perturbative $\alpha^{\prime}$ corrections of string theory. This work is motivated by the Belinskii, Khalatnikov and Lifshitz (BKL) proposal, which suggests that the structure of the black hole interior in vacuum Einstein’s equations can be transformed into the Kasner universe near the singularity. This transformation allows for the description of the black hole interior using the $O\left(d,d\right)$ invariant anisotropic Hohm-Zwiebach action, which includes all orders of $\alpha^{\prime}$ corrections.
@@
Black holes are the most important objects in our universe, as understanding their internal structure and microstate could reveal the fundamental nature of physics. However, gravitational collapse inevitably leads to a curvature singularity at the center of a black hole, a feature that remains unavoidable in the framework of Einstein’s gravity [1, 2]. This suggests that general relativity is incomplete and breaks down in the vicinity of this singularity. String theory, a candidate theory of quantum gravity, is expected to modify the general relativity by including higher-derivative corrections and resolving this longstanding issue. However, to date, it has only provided insights into two-dimensional string black holes [3, 4], leaving unresolved the singularity problem of the four-dimensional black hole in general relativity. This unbridgeable gap arises from the unknown non-perturbative aspects of string theory when approaching the curvature singularity.
@@
where we adopt $116\pi G_{D}=$ 1, $\phi\left(x\right)$ denotes a physical dilaton and we set Kalb-Ramond field $b_{\mu\nu}=0$ for simplicity. Moreover, when $\phi$ is a constant, the action reduces to the Einstein-Hilbert action and we will consider the stringy corrections to this action in this letter. On the other hand, since the spacetime dimension of bosonic string theory is fixed to be $26$, we can see the manifold as a direct product of the black hole and a compact internal space [24]. Due to the fact that the decoupling of the black hole and internal space hold for all order of $\alpha^{\prime}$ corrections [25], we only need to consider the black hole part in the following calculations. To obtain the Kasner universe near the black hole singularity, we first consider the $D=n+3$ dimensional Schwarzschild-Tangherlini black hole [26, 27]:
@@
$ds^{2}=-d\tau^{2}+\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\tau^{2\beta_{i}}dx_{i}^{2}.$ (3)
@@
$\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\beta_{i}^{2}=1,\qquad\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\beta_{i}=1.$ (4)
@@
$\mathcal{S}\equiv\left(\begin{array}[]{cc}0&g_{ij}\\ g_{ij}^{-1}&0\end{array}\right),\quad\mathrm{Tr}\left(\dot{\mathcal{S}}^{2k}\right)=\left(-4\right)^{k}2\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{2k},$ (7)
@@
$I_{\mathrm{HZ}}^{\left(3\right)}=\alpha^{\prime 3}\left(-4\right)^{4}\left[2c_{4,0}\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{8}+4c_{4,1}\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{4}\right)^{2}\right],$ (11)
@@
$\ddot{\Phi}+\frac{1}{2}\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}f_{i}\left(H_{i}\right)$ $=$ $0,$
@@
$\dot{\Phi}^{2}+\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}g_{i}\left(H_{i}\right)$ $=$ $0,$ (12)
@@
$-8\alpha^{\prime 3}4^{4}\left[2c_{4,0}H_{i}^{7}+4c_{4,1}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}H_{j}^{4}\right)H_{i}^{3}\right]+\cdots,$
@@
$-7\alpha^{\prime 3}4^{4}\left[2c_{4,0}H_{i}^{8}+4c_{4,1}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}H_{j}^{4}\right)H_{i}^{4}\right]+\cdots,$
@@
$R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}=\frac{4}{\left(\tau-\tau_{0}\right)^{4}}\left(\frac{1}{2}\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\beta_{i}^{4}-2\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\beta_{i}^{3}+\frac{3}{2}\right).$ (15)
@@
$\ddot{\Omega}-\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}h_{i}\left(H_{i}\right)\right)\Omega$ $=$ $0,$
@@
$\dot{\Omega}^{2}+\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}g_{i}\left(H_{i}\right)\right)\Omega^{2}$ $=$ $0,$ (17)
@@
$H_{i}\left(\tau\right)$ $=$ $\frac{\beta_{i}}{\tau-\tau_{0}}-\frac{4\beta_{i}^{3}+\beta_{i}\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}}{4\left(\tau-\tau_{0}\right)^{3}}\alpha^{\prime}+\cdots,$
$\Omega\left(\tau\right)$ $=$ $\gamma\left(\tau-\tau_{0}\right)+\frac{\gamma}{4\left(\tau-\tau_{0}\right)}\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\beta_{i}^{4}\right)\alpha^{\prime}+\cdots,$
@@
$\Phi\left(\tau\right)=-\log\left(\gamma\left(\tau-\tau_{0}\right)\right)-\frac{1}{4\left(\tau-\tau_{0}\right)^{2}}\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\beta_{i}^{4}\right)\alpha^{\prime}+\cdots.$ (21)
Based on the perturbative solution (LABEL:eq:perturbed_solution) and (21), we obtain the non-perturbative and non-singular solution which satisfies the EOM (12). The dilaton solution is given by
$\Phi\left(\tau\right)=-\frac{1}{2N}\log\left(\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}\right),\qquad\sigma^{2}\equiv 2d\frac{\tau^{2}}{\alpha^{\prime}},$ (22)
where $\lambda_{k}$ is fixed by the coefficients $c_{1}$, … , $c_{k}$, and the corresponding solutions for $f_{i}$, $g_{i}$ and $H_{i}$:
@@
$g_{i}$ $=$ $-\dot{\Phi}^{2}\stackrel{{\scriptstyle[}}{{k}}=1]{\infty}{\sum}\frac{b_{k}}{1+\sigma^{2k-2}},$
$H_{i}$ $=$ $\frac{\sqrt{\alpha^{\prime}}\ddot{\Phi}}{e^{\Phi\left(\tau\right)}\beta_{i}}\stackrel{{\scriptstyle[}}{{k}}=1]{\infty}{\sum}\frac{b_{k}}{1+\sigma^{2k-2}}$ (23)
$-\frac{\sqrt{2d}\dot{\Phi}}{e^{\Phi\left(\tau\right)}\beta_{i}}\stackrel{{\scriptstyle[}}{{k}}=1]{\infty}{\sum}\frac{k\sigma^{2k-1}b_{k+1}}{\left(1+\sigma^{2k}\right)^{2}},$
@@
$\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}b_{1}\left(\beta_{i}\right)=2,\qquad\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}b_{k}\left(\beta_{i}\right)=0,\qquad k>1.$ (24)
@@
$R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ $=$ $2\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{2}\right)^{2}+2\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{4}\right)$ (25)
$+8\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{2}\dot{H}_{i}\right)+4\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\dot{H}_{i}^{2}\right)$
$=$ $\left(\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}\right)^{-\varrho}F\left(\tau\right),$
where $\varrho$ is some positive number and $F\left(\tau\right)$ is a regular function. So the singularities for the spacetime and dilaton appear if and only if
$\mathrm{Singular\;condition}:\qquad\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}=0,$ (26)
@@
$\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}$ $=$ $2099520\sigma^{6}+1283040\sigma^{4}+13680\sigma^{2}$ (27)
@@
This equation has no real roots for $\sigma$, indicating that the singularities of the Schwarzschild and rotating, accreting black holes are eliminated in the non-perturbative solution (see Appendix for details). Indeed, the parameter $\lambda_{k}$ is determined by the condition $c_{n\leq k}$. Considering any order $k$, although $\lambda_{n\leq k}$ are fixed by the coefficients $c_{n\leq k}$, we always have freedom to choose $\lambda_{n>k}$ which violate the singular condition (26), and obtaining the non-singular solutions. Moreover, It is straightforward to verify that this solution (23) covers the perturbative solution (LABEL:eq:perturbed_solution) as $\alpha^{\prime}\rightarrow 0$. Finally, it is worth noting that string effects only resolve the singularity but do not modify the physical properties of black holes beyond the near-singular region.
@@
In conclusion, we demonstrated that the near-singularity geometry of Einstein’s black holes can be effectively described by the Kasner universe. Since, the Kasner universe only depends on $\tau$, we used the Hohm-Zwiebach action to study the singularity problem of black holes. The results indicated that the singularities of Schwarzschild and rotating, accreting black holes may be resolved by considering the $\alpha^{\prime}$ corrections of string theory if the singular condition (26) is broken by the known coefficients $c_{k}$. Nevertheless, we can still make the solution non-singular by adjusting $\lambda_{n>k}$ freely [15]. Moreover, according to the BKL proposal, the general realistic collapse to a spacelike singularity can be described by the Kasner universe, which consists of different Kasner epochs characterized by various choices of the Kasner exponents $\beta_{i}$ [29]. Since our results can apply to any $\beta_{i}$, it suggests that it is possible to study the singularity problem for the astrophysically realistic black holes through string theory.

### v3

@@ Abstract
In this letter, we derive the singular condition for black holes and demonstrate the potential resolution of the Schwarzschild black hole singularity in general relativity using non-perturbative $\alpha^{\prime}$ corrections of string theory. This work is motivated by the Belinskii, Khalatnikov and Lifshitz (BKL) proposal, which suggests that the structure of the black hole interior in vacuum Einstein’s equations can be transformed into the Kasner universe near the singularity. This transformation allows for the description of the black hole interior using the $O\left(d,d\right)$ invariant anisotropic Hohm-Zwiebach action, which includes all orders of $\alpha^{\prime}$ corrections.
@@
Black holes are the most important objects in our universe, as understanding their internal structure and microstate could reveal the fundamental nature of physics. However, gravitational collapse inevitably leads to a curvature singularity at the center of a black hole, a feature that remains unavoidable in the framework of Einstein’s gravity [1, 2]. This suggests that general relativity is incomplete and breaks down in the vicinity of this singularity. String theory, a candidate theory of quantum gravity, is expected to modify the general relativity by including higher-derivative corrections and resolving this longstanding issue. However, to date, it has only provided insights into two-dimensional string black holes [3, 4], leaving unresolved the singularity problem of the four-dimensional black holes of Einstein’s gravity. This unbridgeable gap arises from the unknown non-perturbative aspects of string theory when approaching the curvature singularity.
@@
where we adopt $16\pi G_{D}=1$, $\phi\left(x\right)$ denotes a physical dilaton and we set Kalb-Ramond field $b_{\mu\nu}=0$ for simplicity. Moreover, when $\phi$ is a constant, the action reduces to the Einstein-Hilbert action and we will consider the stringy corrections to this action in this letter. On the other hand, since the spacetime dimension of bosonic string theory is fixed to be $26$, we can see the manifold as a direct product of the black hole and a compact internal space [24]. Due to the fact that the decoupling of the black hole and internal space hold for all order of $\alpha^{\prime}$ corrections [25], we only need to consider the black hole part in the following calculations. To obtain the Kasner universe near the black hole singularity, we first consider the $D=n+3$ dimensional Schwarzschild-Tangherlini black hole [26, 27]:
@@
$ds^{2}=-d\tau^{2}+\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\tau^{2\beta_{i}}dx_{i}^{2}.$ (3)
@@
$\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\beta_{i}^{2}=1,\qquad\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\beta_{i}=1.$ (4)
@@
$\mathcal{S}\equiv\left(\begin{array}[]{cc}0&g_{ij}\\ g_{ij}^{-1}&0\end{array}\right),\quad\mathrm{Tr}\left(\dot{\mathcal{S}}^{2k}\right)=\left(-4\right)^{k}2\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{2k},$ (7)
@@
$I_{\mathrm{HZ}}^{\left(3\right)}=\alpha^{\prime 3}\left(-4\right)^{4}\left[2c_{4,0}\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{8}+4c_{4,1}\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{4}\right)^{2}\right],$ (11)
@@
$\ddot{\Phi}+\frac{1}{2}\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}f_{i}\left(H_{i}\right)$ $=$ $0,$
@@
$\dot{\Phi}^{2}+\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}g_{i}\left(H_{i}\right)$ $=$ $0,$ (12)
@@
$-8\alpha^{\prime 3}4^{4}\left[2c_{4,0}H_{i}^{7}+4c_{4,1}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}H_{j}^{4}\right)H_{i}^{3}\right]+\cdots,$
@@
$-7\alpha^{\prime 3}4^{4}\left[2c_{4,0}H_{i}^{8}+4c_{4,1}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}H_{j}^{4}\right)H_{i}^{4}\right]+\cdots,$
@@
$R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}=\frac{4}{\left(\tau-\tau_{0}\right)^{4}}\left(\frac{1}{2}\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\beta_{i}^{4}-2\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\beta_{i}^{3}+\frac{3}{2}\right).$ (15)
@@
$\ddot{\Omega}-\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}h_{i}\left(H_{i}\right)\right)\Omega$ $=$ $0,$
@@
$\dot{\Omega}^{2}+\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}g_{i}\left(H_{i}\right)\right)\Omega^{2}$ $=$ $0,$ (17)
@@
$H_{i}\left(\tau\right)$ $=$ $\frac{\beta_{i}}{\tau-\tau_{0}}-\frac{4\beta_{i}^{3}+\beta_{i}\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}}{4\left(\tau-\tau_{0}\right)^{3}}\alpha^{\prime}+\cdots,$
$\Omega\left(\tau\right)$ $=$ $\gamma\left(\tau-\tau_{0}\right)+\frac{\gamma}{4\left(\tau-\tau_{0}\right)}\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\beta_{i}^{4}\right)\alpha^{\prime}+\cdots,$
@@
$\Phi\left(\tau\right)=-\log\left(\gamma\left(\tau-\tau_{0}\right)\right)-\frac{1}{4\left(\tau-\tau_{0}\right)^{2}}\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\beta_{i}^{4}\right)\alpha^{\prime}+\cdots.$ (21)
Based on the perturbative solutions (LABEL:eq:perturbed_solution) and (21), we now construct a non-perturbative  [^1] Our statement that this solution is non-perturbative does not rely on the convergence of the series. Rather, it emphasizes that the logarithmic resummation captures contributions that cannot be represented by any finite-order polynomial in $\alpha^{\prime}$. In this sense, the solution extends into the non-perturbative regime, effectively incorporating all orders of $\alpha^{\prime}$ corrections. and non-singular solution that satisfies the EOM (12). Assuming the coefficients $c_{k\leq n}$ are known, the dilaton solution takes the form
$\Phi\left(\tau\right)=-\frac{1}{2N}\log\left(\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}\right),\qquad\sigma^{2}\equiv 2d\frac{\tau^{2}}{\alpha^{\prime}},$ (22)
where $N\geq n-1$ is an arbitrary integer corresponding to the truncation order of the $\alpha^{\prime}$ expansion, and $\lambda_{0}>0$, $\lambda_{n}>0$. The coefficients $\lambda_{N}$, $\lambda_{N-1}$, … , $\lambda_{N-n+1}$ are determined by $c_{1}$, … , $c_{n}$, while the remaining parameters $\lambda_{0}$, $\lambda_{1}$, … , $\lambda_{N-n}$ are free and can be chosen to violate the singularity condition (28), which we will introduce shortly. For example, consider the first two terms of the expansion of (22) in the perturbative regime $\tau\rightarrow\infty$ (or equivalently $\alpha^{\prime}\rightarrow 0$):
$\Phi\left(\frac{\tau}{\alpha^{\prime}}\rightarrow\infty\right)$ $=$ $-\frac{1}{2}\log\left(2d\frac{\tau^{2}}{\alpha^{\prime}}\lambda_{N}^{1/N}\right)$ (23)
$-\frac{1}{2N}\frac{\lambda_{N-1}}{\lambda_{N}}\frac{1}{2d}\frac{\alpha^{\prime}}{\tau^{2}}+\cdots.$
As discussed earlier, the $N$-th order in the $\alpha^{\prime}$ expansion becomes dominant, and the coefficients $\lambda_{N}$ and $\lambda_{N-1}$ can be fixed by matching to the perturbative coefficients $c_{1}$, $c_{2}$. Comparing this expansion with the perturbative solution (21), we determine
$\lambda_{N}$ $=$ $\gamma^{2N}\left(\frac{\alpha^{\prime}}{2d}\right)^{N},$
$\lambda_{N-1}$ $=$ $dN\gamma^{2N}\left(\frac{\alpha^{\prime}}{2d}\right)^{N}\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4},$ (24)
where we have set $\tau_{0}=0$ for simplicity, and used $c_{1}=-1/8$, $c_{2}=1/64$. Higher-order coefficients are listed in the Appendix B. The corresponding functions $f_{i}$, $g_{i}$ and $H_{i}$ are given by
@@
$g_{i}$ $=$ $-\dot{\Phi}^{2}\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{b_{k}}{1+\sigma^{2k-2}},$
$H_{i}$ $=$ $\frac{\sqrt{\alpha^{\prime}}\ddot{\Phi}}{e^{\Phi\left(\tau\right)}\beta_{i}}\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{b_{k}}{1+\sigma^{2k-2}}$ (25)
$-\frac{\sqrt{2d}\dot{\Phi}}{e^{\Phi\left(\tau\right)}\beta_{i}}\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{k\sigma^{2k-1}b_{k+1}}{\left(1+\sigma^{2k}\right)^{2}},$
@@
$\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}b_{1}\left(\beta_{i}\right)=2,\qquad\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}b_{k}\left(\beta_{i}\right)=0,\qquad k>1.$ (26)
@@
$R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ $=$ $2\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{2}\right)^{2}+2\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{4}\right)$ (27)
$+8\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{2}\dot{H}_{i}\right)+4\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\dot{H}_{i}^{2}\right)$
$=$ $\left(\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}\right)^{-\varrho}F\left(\tau\right),$
where $\varrho\equiv 8-\frac{2}{N},$ and $F\left(\tau\right)$ denotes a regular function. The explicit form of $F\left(\tau\right)$ together with a detailed discussion of its possible singular behavior are given in Eq. (43) in Appendix B. So the singularities for the spacetime and dilaton appear if and only if
$\mathrm{Singular\;condition}:\qquad\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}=0,$ (28)
@@
$\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}$ $=$ $2099520\sigma^{6}+1283040\sigma^{4}+13680\sigma^{2}$ (29)
@@
This equation has no real roots for $\sigma$, indicating that the singularities of the Schwarzschild and rotating, accreting black holes are eliminated in the non-perturbative solution (see Appendix for details). Moreover, in this condition, both the string coupling $g_{s}=\exp\left(\phi\right)$ and the curvature invariant $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ in the Einstein frame remain regular. The detailed analysis is provided in Appendix D. Indeed, since our ansatz (22) is constructed under the assumption that the coefficients $c_{k\leq n}$ are known, the parameter $\lambda_{k}$ is determined by the condition $c_{n\leq k}$. Considering any order $k$, although $\lambda_{n\leq k}$ are fixed by the coefficients $c_{n\leq k}$, we always have freedom to choose $\lambda_{n>k}$ which violate the singular condition (28), and obtaining the non-singular solutions. It is worth noting that this ansatz does not represent the unique or complete solution but rather provides a constructive framework to investigate how higher-order $\alpha^{\prime}$ corrections in the $O\left(d,d\right)$-invariant effective action may contribute to singularity resolution. Once the full set of coefficients $c_{k}$ becomes available from string theory (e.g., in heterotic or bosonic models), the solution would need to be correspondingly modified. Moreover, It is straightforward to verify that this solution (25) covers the perturbative solution (LABEL:eq:perturbed_solution) as $\alpha^{\prime}\rightarrow 0$. Finally, it is worth noting that our results focus on the resolution of the singularity. The key point is that the $\alpha^{\prime}$ corrections become significant only when the curvature grows sufficiently large. In the vicinity of the singularity, the inverse gradients of the background fields become comparable to the fundamental string length, such that $\alpha^{\prime}R\geq 1$, $\alpha^{\prime}\left(\nabla\phi\right)^{2}\geq 1$, and so on. In this high-curvature regime, the contributions from all orders of $\alpha^{\prime}$ corrections must be included. Our non-perturbative and non-singular solution is precisely obtained in this region. In the low-curvature regime, where $\alpha^{\prime}\rightarrow 0$, our non-perturbative solution can be expanded perturbatively and reproduces the familiar Schwarzschild black hole of Einstein’s gravity at leading order. Furthermore, we would like to emphasize that there are several indications that the $O\left(d,d\right)$ invariance may also be preserved in the low-curvature black-hole region, although this point is often overlooked in the literature. The Schwarzschild black hole (with vanishing dilaton) is T-dual to the string black hole (with a nontrivial dilaton), as observed in [41, 42, 43]. This duality has recently been successfully incorporated into the $O\left(d,d\right)$-invariant framework at large $D$ in our previous work [44]. It is therefore possible to investigate whether the $O\left(d,d\right)$ invariance can be preserved throughout the entire spacetime and whether the transition between the near-singularity (high-curvature) region and the asymptotic (low-curvature) black-hole region is smooth.
@@
In conclusion, it is believed that the singularity geometry of the Schwarzschild black hole solution in Einstein’s gravity can be effectively described by the Kasner universe. Since the Schwarzschild black hole is a solution of the string low-energy effective action with a constant dilaton, this property allows us to consistently introduce higher-order $\alpha^{\prime}$ corrections, which modify the near-singularity geometry while preserving the Schwarzschild solution at leading order in the low-curvature regime. More specifically, since the Kasner universe only depends on $\tau$, we used the Hohm-Zwiebach action to study the singularity problem of black holes. The results indicated that the singularities of Schwarzschild and rotating, accreting black holes may be resolved by considering the $\alpha^{\prime}$ corrections of string theory if the singular condition (28) is broken by the known coefficients $c_{k}$. Nevertheless, we can still make the solution non-singular by adjusting $\lambda_{n>k}$ freely [15]. Moreover, according to the BKL proposal, the general realistic collapse to a spacelike singularity can be described by the Kasner universe, which consists of different Kasner epochs characterized by various choices of the Kasner exponents $\beta_{i}$ [29]. Since our results can apply to any $\beta_{i}$, it suggests that it is possible to study the singularity problem for the astrophysically realistic black holes through string theory.